Calculate the distance between two objects, 6 000 kg and 6 x 1012 kg respectively, if the magnitude of the 

force between them is 3 x 108 N.



1
Expert's answer
2021-10-25T09:56:32-0400
"F=\dfrac{GMm}{r^2},""r=\sqrt{\dfrac{GMm}{F}},""r=\sqrt{\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times6000\ kg\times6\times10^{12}\ kg}{3\times10^8\ N}},""r=0.09\ m."